3. Two forces of 40.0 N and 30.0 N act on a particle. a) Find the maximum and minimum magnitudes of the resultant of these two forces. b) What is the magnitude of the resultant if the angle between them is 45ᵒ?
1
Expert's answer
2021-10-25T17:55:04-0400

a) The maximum is


"D_\text{max}=30+40=70\text{ N}."

The minimum is


"D_\text{min}=|30-40|=10\text{ N}."

b) The resultant is


"D=\sqrt{30^2+40^2-2·30·40\cos(90°+45°)}=\\
=64.8\text{ N}."